# Generalization of Gibbs and Langevin Monte Carlo Algorithms in the Interpolation Regime

### v1

@@ 3.2 Bounds
$F\left(h,\mathbf{x}\right)\leq\Gamma\left(h,\mathbf{x},\bm{\beta}\right)+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]+\ln\left(1/\delta\right),$
@@
$\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]\leq\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\right]+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]+\ln\left(1/\delta\right).$
@@
To prove (i), we apply this to the random variable $Y=F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$ on the probability space $\left(\mathcal{H\times X}^{n},\Omega\otimes\Sigma^{\otimes n},\rho_{G_{\beta}}\right)$ as defined in (2). Together with the definition of the Gibbs posterior, this gives, with probability at least $1-\delta$ in $\left(h,\mathbf{x}\right)\sim\rho_{G_{\beta}}$ (equivalent to saying $\mathbf{x}\sim\mu^{n}$ and $h\sim G_{\beta}\left(\mathbf{x}\right)$),
$\leq\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim G_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(g,\mathbf{x}\right)+\beta\hat{L}\left(g,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+\ln\left(1/\delta\right)$
$=\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)+\beta\hat{L}\left(g,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)-\beta\hat{L}\left(g,\mathbf{x}\right)-\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+\ln\left(1/\delta\right)$
@@
Subtract $\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$ and use (4). For (ii) apply Markov’s inequality to $\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]$ instead. By Jensen’s inequality
$e^{\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]}\leq\mathbb{E}_{g\sim G_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(g,\mathbf{x}\right)+\beta\hat{L}\left(g,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]$
and proceed as above using (4). ∎
Up to the application of (4), the above proof of (i) just gives the single-draw version of the PAC-Bayesian bound as in Rivasplata et al. (2020) applied to the Gibbs posterior, while (ii) is the standard PAC-Bayesian bound applied to the Gibbs posterior until (4) is invoked.
3.3 Loss functions and secondary loss functions
To apply Theorem 3.2, we need to control the exponential moment $\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]$, but otherwise we have free choice of the function $F$. This gives the method some flexibility. If $\exp\left(F\right)$ has sufficient integrability properties, we can exchange the two expectations and often there is a bound on $\mathbb{E}_{\mathbf{x}}\left[e^{F\left(g,\mathbf{x}\right)}\right]$ uniform in $g$, which then carries over to $\mathbb{E}_{g\sim\pi}\mathbb{E}_{\mathbf{x}}\left[e^{F\left(g,\mathbf{x}\right)}\right]$, since $\pi$ is a probability measure. In this way bounds for sub-Gaussian or sub-exponential losses can be obtained, but also for U-statistics or even non-iid data, sampled from the trajectories of time-homogeneous, ergodic Markov chains. In Section B.1 in the appendix, we derive a bound for sub-Gaussian losses from Theorem 3.2; other examples are planned for a longer version of the paper.
The function $F$ may be defined in terms of other, application-dependent loss functions, which are different from the loss $\ell$, which defines the Gibbs posterior and the functional $\Gamma$. To illustrate this point, let $f:\mathcal{H}\times\mathcal{X}\rightarrow[0,1]$ be measurable and set $F(h,\mathbf{x})=n~\kappa\left(\frac{1}{n}\sum_{i}f(h,x_{i}),\mathbb{E}_{x}\left[f(h,x)\right]\right)$, with $\kappa$ the relative entropy as in (1). Then, Theorem 1 of Maurer (2004) gives $\mathbb{E}_{\mathbf{x}}\left[e^{F(h,\mathbf{x})}\right]\leq 2\sqrt{n}$ for $n\geq 8$. Substitution in Theorem 3.2 and division by $n$ then give the following corollary.
Corollary 3.3.
Let $f:\mathcal{H}\times\mathcal{X}\rightarrow[0,1]$ be measurable, $\delta>0$ and $n\geq 8$. Then, with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$ and $h\sim G_{\beta}(\mathbf{x})$

### v2

@@ 4.2 Stability of the Bounds
$F\left(h,\mathbf{x}\right)$ $\leq$ $-\beta\hat{L}\left(h,\mathbf{x}\right)+\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
$+R_{\infty}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)+\Delta.$
If $F$ and $\ell$ are bounded, then $R_{\infty}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)$ can be replaced by
$\max\left\{0,\beta\left\|\ell\right\|_{\infty}+\left\|F\right\|_{\infty}+\ln\sqrt{2KL\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)}\right\}.$
(ii) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$
$\mathbb{E}_{\nu\left(\mathbf{x}\right)}\left[F(h,\mathbf{x})\right]$ $\leq$ $-\beta\ \mathbb{E}_{\nu\left(\mathbf{x}\right)}\big[\hat{L}\left(h,\mathbf{x}\right)\big]+\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
$+KL\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)+\Delta.$
The left-hand side in both inequalities is the random variable that we wish to bound, depending on the distribution $\nu$. The right-hand side of the top line is the bound, as computed from the distributions $\nu(\mathbf{x})$ and $\nu_{0}^{K-1}(\mathbf{x})$ and includes the dependence on the exponential moment of $F$ and the confidence parameter $\delta$. The second line gives the error incurred by the fact that none of the distributions is really the right Gibbs distribution. The first term there gives the error for the target distribution $\nu(\mathbf{x})$, and is different in the single-draw and classical PAC-Bayesian cases. The term $\Delta$ results from approximating the temperature integral by the expectations in a finite number of distributions, as described in Lemma 4.4.
The amendment to (i) is necessary, since we know of no process with useful bounds on $R_{\infty}$. The replacement indeed converges to zero with relative entropy, but, since $\beta\left\|\ell\right\|_{\infty}+\left\|F\right\|_{\infty}$ is typically of order $n$, it requires the relative entropy to be exponentially small in $n$. Nevertheless, if the bounds in Corollary C.2 are substituted in (ii) or in the amendment to (i), they guarantee, with sufficient computational budget and appropriate choices of $t$ and $\eta$, the convergence of LMC to our bounds for the Gibbs posterior.
A detailed proof of Theorem 4.5 is given in Appendix C.4. Part (ii) and (i) without amendment follow more or less directly from the PAC-Bayesian theorem, Lemma 4.4 and equation (6) and a reasoning using $R_{\infty}$ analogous to (6). The amendment to (i) requires a special adapted proof of the PAC-Bayesian theorem.
5 Experiments
The purpose of our experiments is twofold. For one, we want to verify that the theoretical dependence of the generalization performance of the Gibbs posterior at low temperatures on its training errors at high temperatures carries over to practical temperature-regularized algorithms like SGLD in real world settings, with overparametrized neural networks. Indeed, in all our experiments the temperature plots of the mean training errors, computed as described below, verify the qualitative prediction that both the failure of generalization for random labels and the success for true labels are related to the areas under the curves at higher temperatures (see for example Figure 2).
Second, we would like to use the training data, and only the training data, to make realistic quantitative predictions of test errors in such settings. This is more difficult, since the guarantees of Corollary C.2 in combination with Theorem 4.5 are inadequate in high-dimensional situations. We achieve this goal with a principled calibration scheme described below.
5.1 Experimental Environment and Algorithms
The real-world data are either the MNIST dataset, subdivided into the two classes of characters 0–4 and 5–9, or the CIFAR-10 dataset to distinguish between animals and vehicles. For impossible data, we randomize the labels of the training data. Our experiments are computationally heavy, so we generally use small sample sizes, from 2000 to 8000 examples. The hypothesis space is the set of weight vectors for a neural network with ReLU activation functions constrained by a Gaussian prior distribution with $\sigma=5$. Neural network architectures are described in Section D.1.1 of the appendix. To approximately sample the weight vectors in the vicinity of the Gibbs posterior, we use ULA as in (15) or SGLD (Welling and Teh, 2011) with constant step size $\eta$.
5.2 The Loss Function $\ell$
Most experiments were done with bounded loss functions $\ell$, either bounded binary cross-entropy as described in Appendix D of Dziugaite and Roy (2018) or the Savage loss (Masnadi-Shirazi and Vasconcelos, 2008). As unbounded loss function we tried binary cross-entropy (BCE) (Section D.2.6), but with a smaller value of $\sigma$, so as to avoid excessive training errors for small values of $\beta$. We compute bounds for the 0-1 loss, using the method described in Section B.2.
5.3 Approximating the Ergodic Mean
As we know of no sufficient criterion for convergence, we terminate iterations at time $T$, when a very slow running mean $\mathbb{M}_{\rm stop}$ of the loss trajectory $\big(\hat{L}(h_{\beta_{k},t},\mathbf{x})\big)_{t=0}^{T}$ stops decreasing. A second running mean $\mathbb{M}_{\rm erg}$ is used as an approximation of the ergodic mean and thus of expectations in the invariant distribution. We thus replace all expectations $\mathbb{E}_{h\sim G_{\beta_{k}}}\big[\hat{L}(h,\mathbf{x})\big]$ occurring in the bounds by $\mathbb{M}_{\rm erg}\big[\big(\hat{L}(h_{\beta_{k},t},\mathbf{x})\big)_{t=0}^{T}\big]$. Both running means $\mathbb{M}_{\rm stop}$ and $\mathbb{M}_{\rm erg}$ are implemented as first-order, recursive lowpass filters described in Section D.1.3 of the appendix.
5.4 Computation of the Bounds and Calibration
For the 01-error (cf. Section B.2) we compute our bounds from the PAC-Bayesian theorem in the form
$\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[L_{\text{0}1}\left(h\right)\right]\leq\kappa^{-1}\left(\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[\hat{L}_{\text{0}1}\left(h,\mathbf{x}\right)\right],\frac{\mathcal{Q}+\ln\left(\frac{2\sqrt{n}}{\delta}\right)}{n}\right),$ (7)
where $\nu_{\beta}\left(\mathbf{x}\right)$ is an LMC approximation to $G_{\beta}\left(\mathbf{x}\right)$ and $\mathcal{Q}$ is a proxy for $KL\left(\nu_{\beta}\left(\mathbf{x}\right),\pi\right)$, the computation of which is described below. Using $\kappa^{-1}$ is somewhat more accurate than the analog of (2). The function $\kappa$ is the relative entropy of two Bernoulli variables, the derivation of the above bound and the inverse function $\kappa^{-1}$ are explained in section B.2.
Our approximation to $KL\left(\nu_{\beta}\left(\mathbf{x}\right),\pi\right)$ is
$A=-\beta\ \mathbb{E}_{h\sim\nu_{\beta}}\left[\hat{L}\left(h,\mathbf{x}\right)\right]+\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K}).$
For a rigorous bound following Theorem 4.5 we would have to set $\mathcal{Q}=A$ plus all the terms bounding the approximation errors as in Corollary C.2. Unfortunately, the quantities $R$ and $\alpha$ are impossible to estimate in practice. But even if we assume these to be in the order of unity, the bounds are too coarse to distinguish between different temperatures with realistic step sizes. A simple calculation (see Lemma C.7 in Appendix C.5) shows that
$\text{KL}\left(G_{\beta},G_{2\beta}\right)\leq\beta\left(\mathbb{E}_{G_{\beta}}\left[\hat{L}\right]-\mathbb{E}_{G_{2\beta}}\left[\hat{L}\right]\right)\leq\beta\mathbb{E}_{G_{\beta}}\left[\hat{L}\right$].
By Corollary C.2 we should therefore have at least $8\eta dR^{2}/\alpha<\mathbb{E}_{G_{\beta}\left(\mathbf{x}\right)}\left[\hat{L}\right$] to distinguish between the expectations in the Gibbs posterior for $\beta$ and $2\beta$. The smallest neural network we use has $d=392,500$. If $\ell$ has values in $\left[0,1\right]$ then $\mathbb{E}_{G_{\beta}\left(\mathbf{x}\right)}\left[\hat{L}\right]\leq 1$, so even if $R$ and $\alpha$ are set to $1$, we would need step sizes in the order of $10^{-7}$. Safe values of $\eta$, as suggested by the theoretical results in Section 4, are therefore impossible in practice, and the bound has to be adapted to a realistic choice of $\eta$.
But the data for the true labels $\mathbf{x}$ and the data $\mathbf{\bar{x}}$ for the random labels are strongly related. Dimension and input marginals are equal. This suggests that the inaccuracy of the LMC approximations affects both in a similar way, and we make the following calibration assumption:
$\frac{KL\left(\nu_{\beta}\left(\mathbf{x}\right),\pi\right)}{KL\left(\nu_{\beta}\left(\mathbf{\bar{x}}\right),\pi\right)}=\frac{A}{\bar{A}},$
where $\bar{A}$ is given by the expression analogous to $A$ for the random labels, compare Figure 2.
The expected error of the random labels in binary classification is $1/2.$ Now let $r$ be the smallest positive factor such that $r\bar{A}$, when substituted for $\mathcal{Q}$ in (7) yields a value greater or equal $1/2$. If the PAC-Bayesian theorem is tight, then this means that $r\bar{A}\geq KL\left(\nu_{\beta}\left(\mathbf{\bar{x}}\right),\pi\right)$. Our calibration assumption above then also implies that
$rA\geq KL\left(\nu_{\beta}\left(\mathbf{x}\right),\pi\right),$
and substitution of $rA$ for $\mathcal{Q}$ in (7) should result in a correct upper bound for the true labels. For a precise definition of $r$ see Section C.6.
Clearly there is no rigorous justification of this argument. It is an experimental finding that our choice of $r$ leads to correct and surprisingly tight upper bounds on the test error of correctly labeled data in all cases we tried. We emphasize that our calibration procedure depends only on the training data.
5.5 Results
Figure 4: SGLD on MNIST and CIFAR-10 with 8000 training examples, MNIST above and CIFAR-10 below, random labels on the left, correct labels on the right. Both random and true labels are trained with the same algorithm and parameters on a fully connected ReLU network with two hidden layers of 1000 and 1500 units, respectively. The calibration factor for MNIST is 0.77, for CIFAR-10 0.89. Train error, test error and our bound for the Gibbs posterior average of the 0-1 loss are plotted against $\beta$.
Several experiments confirm the validity of the proposed bounds. An example is shown in Figure 4, where a fully connected ReLU network with two hidden layers of 1000 (respectively 1500) units each is trained with SGLD at inverse temperatures $\beta=$ 0, 500, 1000, 2000, 4000, 8000, 16000, 32000, and 64000. The train error for random labels is about 0.1 (or even less) at $\beta=64000$, where the bound is above 0.5. The test error for correct labels, however, is tightly bounded above.
Notice that for MNIST, which has the tightest bounds, the training error for the true labels is rapidly decreasing from 0.5 to 0.17 at $\beta=500$ and to 0.1 at $\beta=1000$. The more moderate initial decrease for CIFAR-10 corresponds to the tendency to overfit on this more difficult dataset. This confirms the intuition, that good generalization at low temperatures is already announced in the high-temperature regime. Experimental bounds for single posterior draws, along with additional experiments including applications to model selection, are presented in Section D.2.
6 Conclusion
Using the integral representation of the log-partition function, the Gibbs posterior admits the computation of upper bounds on the true error based on the training data and for any temperature. These bounds are stable under perturbation in relative entropy and can be approximated by Langevin Monte Carlo (LMC) algorithms. However, for realistic experiments, the approximations obtained by these algorithms are coarse and require calibration, which leads to rather tight bounds in the interpolation regime of overparametrized neural networks.
Appendix
In this appendix, we provide a glossary of notation, give additional theoretical results and missing proofs, and provide more information on the numerical experiments, as well as additional experimental results.
Appendix A Table of notation
Notation Brief description Section
$\mathcal{X}$ space of data 2, 3, 4
$\Sigma$ sigma algebra (events) on $\mathcal{X}$ 2
$\mu$ probability of data 2, 4.2
$n$ sample size 1, 2, 3, 4
$\delta$ confidence parameter in high probability bounds 1, 2, 4.2
$\mathbf{x}$ generic member $\left(x_{1},...,x_{n}\right)\in\mathcal{X}^{n}$, training sample 1, 2, 3, 4
$\mathcal{H}$ hypothesis space 2, 3, 4.2
$\Omega$ sigma algebra (events) on $\mathcal{H}$ 2, 3
$\ell$ $\ell:\mathcal{H\times X\rightarrow}\left[0,\infty\right)$ loss function 1, 2, 4.2, 5
$\mathcal{P}\left(\mathcal{H}\right)$ probability measures on $\mathcal{H}$ 2, 3
$\pi$ nonnegative a-priori measure on $\mathcal{H}$ 2, 3, 4
$\|f\|_{\infty}$ The sup-norm is defined: $\|f\|_{\infty}=\sup\{|f(s)|:s\in S\}$ 2, 4
$\sigma$ width of Gaussian prior 4.1, 5
$L\left(h\right)$ $L\left(h\right)=\mathbb{E}_{x\sim\mu}\left[\ell(h,x)\right]$, expected loss of $h\in\mathcal{H}$ 2
$\hat{L}\left(h,\mathbf{x}\right)$ $\hat{L}\left(h,\mathbf{x}\right)=\left(1/n\right)\sum_{i=1}^{n}\ell\left(h,x_{i}\right)$, empirical loss of $h\in\mathcal{H}$ 2, 3, 4, 5
$\beta$ inverse temperature 1, 2, 3, 4, 5
$Z_{\beta}\left(\mathbf{x}\right)$ partition function 3, 4.1
$G_{\beta,\pi}\left(\mathbf{x}\right)$ Gibbs posterior with energy $\hat{L}$ and prior $\pi$ 1, 3, 4, 5
$\mathbb{E}_{g\sim G_{\beta}\left(\mathbf{x}\right)}$ posterior expectation 3
$\beta_{0}^{K}$ increasing sequence $\left(0=\beta_{0}<\beta_{1}<\cdots<\beta_{K}=\beta\right)$ of positive reals 4.2, 5
$\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})$ bounding functional 4.2, 5
$F\left(h,\mathbf{x}\right)$ placeholder for generalization gap 2, 4.2
$\kappa(p,q)$ $kl\left(p,q\right)=p\ln\frac{p}{q}+\left(1-p\right)\ln\frac{1-p}{1-q}$, rel. entropy of Bernoulli variables 5
$\kappa^{-1}(p,t)$ $\inf\left\{q:q\geq p,\kappa\left(p,q\right)\geq t\right\}$ 5.4
$\text{KL}\left(\rho,\nu\right)$ $\int\left(\ln\frac{d\rho}{d\nu}\right)d\rho$, KL-divergence of $\rho,\nu\in\mathcal{P}\left(\mathcal{H}\right)$ 2, 3, 4, 5
$R_{\infty}\left(\rho,\nu\right)$ $\sup_{h\in\mathcal{H}}\ln\frac{d\rho}{d\nu}(h)$, Rényi-infinity divergence of $\rho,\nu\in\mathcal{P}\left(\mathcal{H}\right)$ 2, 4.2
$d_{TV}\left(\rho,\nu\right)$ total variation distance C.4
$W_{p}\left(\rho,\nu\right)$ $p$-Wasserstein distance C.4
$\eta$ step size or learning rate C.1
$\nu_{\beta,\eta}$ invariant measure of LMC approximation of $G_{\beta}$ with step size $\eta$ C.1
$\nu_{\beta,\eta,t}$ LMC approximation of $G_{\beta}$ with step size $\eta$ at iteration $t$ C.1
$\nu_{\beta,\epsilon}$ invariant measure of CLD process with step size $\epsilon$ C.1
$\nu_{\beta,\epsilon,t}$ CLD with step size $\epsilon$ at iteration $t$ C.1
$r$ calibration factor 5.4
$\mathbf{\bar{x}}$ randomly labeled data 5.4
$\mathbb{M}_{\rm stop}$$\mathbb{M}_{\rm erg}$ filters for stopping and ergodic mean 5.3
$L_{01}\left(h\right)$ $L_{01}\left(h\right)=\mathbb{E}_{x\sim\mu}\left[\ell_{01}(h,x)\right]$, expected 01-loss of $h\in\mathcal{H}$ 5.4
$\hat{L}_{01}\left(h,\mathbf{x}\right)$ $\hat{L}\left(h,\mathbf{x}\right)=\left(1/n\right)\sum_{i=1}^{n}\ell_{01}\left(h,x_{i}\right)$, empirical 01-loss of $h\in\mathcal{H}$ 5.4
Appendix B PAC-Bayes
We review PAC-Bayesian theory. There is no claim to novelty.
B.1 Proof of the PAC-Bayesian Theorem
Definition B.1.
Then, $\mathbb{E}_{\left(h,\mathbf{x}\right)\sim\rho_{\nu}}\left[\phi\left(h,\mathbf{x}\right)\right]=\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[\phi\left(h,\mathbf{x}\right)\right]$ for measurable $\phi:\mathcal{H}\times\mathcal{X}^{n}\rightarrow\mathbb{R}$. To draw the pair $\left(h,\mathbf{x}\right)$ from $\rho_{\nu}$ we first draw the training sample $\mathbf{x}$, and then sample $h$ from $\nu\left(\mathbf{x}\right)$.
Restatement of Theorem 2.1
Theorem B.2.
[…]
(i) for $\delta>0$ with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu\left(\mathbf{x}\right)$
$F\left(h,\mathbf{x}\right)\leq\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}\left(h\right)+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]+\ln\left(\frac{1}{\delta}\right)$
@@ B.1 Proof of the PAC-Bayesian Theorem
$\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]\leq\text{KL}\left(\nu\left(\mathbf{x}\right),\pi\right)+\ln\frac{\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]}{\delta}$
@@
To prove (i), we apply this to the random variable $Y=F\left(h,\mathbf{x}\right)-\ln\left(d\left(\nu\left(\mathbf{x}\right)/d\pi\right)\right)\left(h\right)$ on the probability space $\left(\mathcal{H\times X}^{n},\Omega\otimes\Sigma^{\otimes n},\rho_{\nu}\right)$. This gives, with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu\left(\mathbf{x}\right)$),
$F\left(h,\mathbf{x}\right)-\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}\left(h\right)$
$\leq\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\nu\left(\mathbf{x}\right)}\left[e^{F\left(g,\mathbf{x}\right)-\ln\left(d\left(\nu\left(\mathbf{x}\right)/d\pi\right)\right)\left(g\right)}\right]+\ln\left(1/\delta\right)$
$=\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)-\ln\left(d\left(\nu\left(\mathbf{x}\right)/d\pi\right)\right)\left(g\right)+\ln\left(d\left(\nu\left(\mathbf{x}\right)/d\pi\right)\right)\left(g\right)}\right]+\ln\left(1/\delta\right)$
@@
For (ii) apply Markov’s inequality on the probability space $\left(\mathcal{X}^{n},\Sigma^{\otimes n},\mu^{n}\right)$ to the random variable $Y=\mathbb{E}_{g\sim\nu\left(\mathbf{x}\right)}\left[F\left(g,\mathbf{x}\right)\right]-KL\left(\nu\left(\mathbf{x}\right),\pi\right)=\mathbb{E}_{g\sim\nu\left(\mathbf{x}\right)}\left[F\left(g,\mathbf{x}\right)-\ln\left(d\left(\nu\left(\mathbf{x}\right)/d\pi\right)\right)\left(g\right)\right]$ instead. By Jensen’s inequality
$e^{\mathbb{E}_{g\sim\nu\left(\mathbf{x}\right)}\left[F\left(g,\mathbf{x}\right)-\ln\left(d\left(\nu\left(\mathbf{x}\right)/d\pi\right)\right)\left(g\right)\right]}\leq\mathbb{E}_{g\sim\nu\left(\mathbf{x}\right)}\left[e^{F\left(g,\mathbf{x}\right)-\ln\left(d\left(\nu\left(\mathbf{x}\right)/d\pi\right)\right)\left(g\right)}\right].$
Then proceed as before. ∎
B.2 Concrete PAC-Bayesian Bounds and the 0-1 Loss
Assume that $\ell$ has values in $[0,1]$. To derive (2) from Theorem 2.1 let $F\left(h,\mathbf{x}\right)=n~\kappa\left(\hat{L}\left(h,\mathbf{x}\right),L\left(h\right)\right)$, where $\kappa$ is the relative entropy of two Bernoulli variables with expectations $p$ and $q$
for which Tolstikhin and Seldin (2013) give the inversion rule $\kappa\left(p,q\right)\leq B$ $\implies$ $q-p\leq\sqrt{2pB}+2B$. Tonelli’s theorem and Theorem 1 of Maurer (2004) then give $\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]=\mathbb{E}_{g\sim\pi}\mathbb{E}_{\mathbf{x}}\left[e^{F\left(g,\mathbf{x}\right)}\right]\leq 2\sqrt{n}$ for $n\geq 8$. Substitution in (i) of Theorem 2.1 and division by $n$ then gives with high probability as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu\left(\mathbf{x}\right)$
$\kappa\left(\hat{L}\left(h,\mathbf{x}\right),L\left(h\right)\right)\leq\frac{1}{n}\left(\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right),$ (10)
and the inversion rule of Tolstikhin and Seldin (2013) gives (2). The derivation of the bound for $\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[L\left(h\right)\right]$ from Theorem 2.1 (ii) is analogous, we just have to use $\kappa\left(\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right],\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[L\left(h\right)\right]\right)\leq\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\kappa\left(\hat{L}\left(h,\mathbf{x}\right),L\left(h\right)\right)$ by the joint convexity of $\kappa$ (see Cover (1999)). This gives